# Unraveling codes: fast, robust, beyond-bound error correction for DRAM

### v1

@@ I Introduction
To achieve CC, a designer sets $t$ to the number of symbols sent by each device, necessitating two extra devices to store the $2t$ parity symbols. This principle is used in typical configurations of both DDR4 and DDR5 server memory [16, 17]. In DDR4 each operation accesses 18 devices, exchanging 4 bytes (4B) per device for a total of 72B. In DDR5 each operation accesses 10 devices, exchanging 8B per device for a total of 80B. Typically the user data is a 64B cache line in both cases, leaving 8B or 16B for parity respectively, and in both cases allowing for a full device to be corrected within-bound. But more recently, e.g., in the Intel Skylake and Cascade Lake server generations [24], some of the extra bits are used as metadata for other purposes than error correction: to store ownership, security information, and cache coherency state; to mark failed areas of DRAM, etc. This leaves less room for parity, so that beyond-bound Chipkill-correct (BBCC) is required. BBCC cannot be perfectly reliable, but it can correct the vast majority of single-device errors.
The goal of a more powerful decoding algorithm must be balanced against implementation constraints. The implementation needs to achieve low-latency encoding and decoding [^2] Because memory errors are infrequent, it is possible to use a decoder which detects errors quickly but is slow to correct them. However, this adds a complexity cost, since such a decoder can drop transactions and/or return them out of order. Some systems must also maintain good performance even when one device develops a persistent fault., small area, and low power consumption. Beyond-bound decoding is more complex than within-bound decoding, and some approaches rely in part on exhaustive search when errors cannot be decoded analytically [4, 14, 36]. The complexity can be reduced, and the reliability improved, by limiting beyond-bound correction to single-device errors only, while performing multiple-device correction within-bound (e.g. up to all 4 DQ of a single device, or 3 DQ across multiple devices). In higher-reliability systems, beyond-bound double-device correction may be required; in that case, exhaustive search is quadratically more expensive.
@@ II-D Motivating application: DDR5 RAM, CC and DCC
ptBurst length 16 bits10 devicesdeviceMix
@@
More conservative systems require even more powerful error correction, such as double-Chipkill-correct (DCC) which can tolerate the failure of more than one device. This can be achieved, for example, using a DDR5 module in lockstep mode, where 128B of user data and 32B of parity are stored across 20 devices instead of 10. Beyond-bound DCC is more complex: for example, exhaustively searching for double-device failures would require trying $\binom{20}{2}=190$ combinations.
@@ II-E Our contribution: Unraveling Reed-Solomon codes
RS over $\mathbb{F}_{2^{16}}$ ${\color[rgb]{0,0.45,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.45,0}\pgfsys@color@rgb@stroke{0}{0.45}{0}\pgfsys@color@rgb@fill{0}{0.45}{0}\leq 15}$ 3 6 ${\color[rgb]{0,0.45,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.45,0}\pgfsys@color@rgb@stroke{0}{0.45}{0}\pgfsys@color@rgb@fill{0}{0.45}{0}3.61\mathrm{e}{-14}}$ $3.61\mathrm{e}{-14}$ no $3.20\mathrm{e}{-14}$ $3.20\mathrm{e}{-14}$ $\leq 7$ 174%
IRS $\ell=2$ 8 3 6 ${\color[rgb]{0,0.45,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.45,0}\pgfsys@color@rgb@stroke{0}{0.45}{0}\pgfsys@color@rgb@fill{0}{0.45}{0}3.61\mathrm{e}{-14}}$ $1.44\mathrm{e}{-7}$ yes $3.20\mathrm{e}{-14}$ $4.66\mathrm{e}{-10}$ 4 106%
@@
f  URS $\ell=8$ 15 1 CC ${\color[rgb]{0,0.45,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.45,0}\pgfsys@color@rgb@stroke{0}{0.45}{0}\pgfsys@color@rgb@fill{0}{0.45}{0}3.55\mathrm{e}{-14}}$ ${\color[rgb]{0,0.45,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.45,0}\pgfsys@color@rgb@stroke{0}{0.45}{0}\pgfsys@color@rgb@fill{0}{0.45}{0}3.55\mathrm{e}{-14}}$ yes $3.55\mathrm{e}{-15}$ $3.55\mathrm{e}{-15}$ 7 34%
@@ IV Application in a DDR5 memory controller
8 $(80,65)$ $(40,32)\!\times\!(40,33)$ 3 $(10,8)^{7}\!\times\!(10,9)$ $1.39\mathrm{e}{-17}$ 8 $1.41\mathrm{e}{-16}$
@@
We chose unraveling codes because of their very low SDC rate, similar to 16-bit codes but requiring only 8-bit field arithmetic; and to enable BBCC without exhaustive search.
@@ V Comparison to Other Error Correction Techniques
V-A Comparison to other large-codeword techniques
@@
DUO ECC goes further, by repurposing the on-die ECC data to provide better controller-level error correction with a long codeword, at the cost of slightly longer transfers, possibly enabling probabilistic CC or even DCC for DIMM configurations that would otherwise not support them. For example, DUO suggests DDR4 ×$\times 4$ and DDR4 ×$\times 4$ modes with 18 and 9 devices, 34 and 68 bits per device, respectively, both with a total of $34\cdot 18=612$-bit codewords. These use a $\mathrm{RS}(2^{8};76,64)$ code which can correct up to 6 byte errors. Or they can probabilistically correct a single-device error for DDR5, or even a double-device error for DDR4, by exhaustive search: that is, by attempting decoding with each device (or pair of devices) erased.
As with Bamboo codes, URS codes can complement DUO, but not as effectively because there are not a whole number of symbols per device, which necessitates shortening the codes [^7] Unless 17-bit symbols are used, but this is worse unless significant metadata is also desired, because it would have a $31\cdot 17=527$-bit payload.. For example, we could use $\mathbb{F}_{2^{9}}$ and divide the 100 extra bits into 1 bit of metadata and 11 nine-bit parity symbols. This would give an $\mathrm{RS}(2^{9};72,61)$ code with 4 or 8 symbols per device for DDR4 or DDR5, shortened from 72 or 36 bits to 68 or 34 bits per device, respectively. In DDR4 this corrects a single-device plus one error efficiently and reliably; a double-device error with failure probability about $2^{-18}\approx 3.81\mathrm{e}{-6}$; or a single-device plus two-symbol error with failure probability about $2^{-9}\approx 0.2\%$. With DDR5, it can correct a single-device failure except with probability about $2^{-27}\approx 3.35\mathrm{e}{-8}$. These calculations are all based on analytically solving the key equation rather than iterating over each device. We could fall back to the iterative solution if the analytic solution fails, which would be a nearly identical solution to DUO but with slightly higher efficiency due to the sparse erasure locator polynomial.
@@ VI Conclusion and future work
VI-A Intellectual property disclosure

### v2

@@ I Introduction
To achieve CC, a designer sets $t$ to the number of symbols sent by each device, necessitating two extra devices to store the $2t$ parity symbols. This principle is used in typical configurations of both DDR4 and DDR5 server memory [16, 17]. In DDR4 each operation accesses 18 devices, exchanging 4 bytes (4B) per device for a total of 72B. In DDR5 each operation accesses 10 devices, exchanging 8B per device for a total of 80B. Typically the user data is a 64B cache line in both cases, leaving 8B or 16B for parity respectively, and in both cases allowing for a full device to be corrected within-bound. But more recently, e.g., in the Intel Skylake and Cascade Lake server generations [25], some of the extra bits are used as metadata for other purposes than error correction: to store ownership, security information, and cache coherency state; to mark failed areas of DRAM, etc. This leaves less room for parity, so that beyond-bound Chipkill-correct (BBCC) is required [^2] Alternatively, extra parity symbols could be written elsewhere [39, 18], but this costs performance and energy.. BBCC cannot be perfectly reliable, but it can correct the vast majority of single-device errors.
The goal of a more powerful decoding algorithm must be balanced against implementation constraints. The implementation needs to achieve low-latency encoding and decoding [^3] Because memory errors are infrequent, it is possible to use a decoder which detects errors quickly but is slow to correct them. However, this adds a complexity cost, since such a decoder can drop transactions and/or return them out of order. Some systems must also maintain good performance even when one device develops a persistent fault., small area, and low power consumption. Beyond-bound decoding is more complex than within-bound decoding, and some approaches rely in part on exhaustive search when errors cannot be decoded analytically [14, 37]. Such a system might correct only single-device errors, or might combine this and within-bound correction of multi-device errors (e.g. up to all 4 DQ of a single device, or 3 DQ across multiple devices). Or, by limiting a search to single-device errors only, the search only needs one computation per device, which may be done in parallel [4]. In higher-reliability systems, beyond-bound double-device correction may be required; in that case, searching for failed devices is quadratically more expensive.
@@ II-D Motivating application: DDR5 RAM, CC and DCC
Burst length 16 bits10 devicesdeviceMix
@@
More conservative systems require even more powerful error correction, such as double-Chipkill-correct (DCC) which can tolerate the failure of more than one device. This can be achieved, for example, using a DDR5 module in lockstep mode, where 128B of user data and 32B of parity are stored across 20 devices instead of 10. Beyond-bound DCC is more complex: for example, searching through all double-device failures would require trying $\binom{20}{2}=190$ combinations.
@@ II-E Our contribution: Unraveling Reed-Solomon codes
RS over $\mathbb{F}_{2^{16}}$ ${\color[rgb]{0,0.45,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.45,0}\leq 15}$ 3 6 ${\color[rgb]{0,0.45,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.45,0}3.61\mathrm{e}{-14}}$ $3.61\mathrm{e}{-14}$ no $3.20\mathrm{e}{-14}$ $3.20\mathrm{e}{-14}$ $\leq 7$ 174%
IRS $\ell=2$ 8 3 6 ${\color[rgb]{0,0.45,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.45,0}3.61\mathrm{e}{-14}}$ $1.44\mathrm{e}{-7}$ yes $3.20\mathrm{e}{-14}$ $4.66\mathrm{e}{-10}$ 4 106%
@@
f  URS $\ell=8$ 15 1 CC ${\color[rgb]{0,0.45,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.45,0}3.55\mathrm{e}{-14}}$ ${\color[rgb]{0,0.45,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.45,0}3.55\mathrm{e}{-14}}$ yes $3.55\mathrm{e}{-15}$ $3.55\mathrm{e}{-15}$ 7 34%
@@ IV Application in a DDR5 memory controller
8 $(80,65)$ $(40,32)\!\times\!(40,33)$ 3 $(10,8)^{7}\!\times\!(10,9)^{\phantom{2}}$ $1.39\mathrm{e}{-17}$ 8 $1.41\mathrm{e}{-16}$
@@
We chose unraveling codes because of their very low SDC rate, similar to 16-bit codes but requiring only 8-bit field arithmetic; and to enable efficient BBCC.
@@ V Comparison to Other Error Correction Techniques
V-A Comparison to MUSE-ECC
In this work we have been assuming a symmetric channel, but in real DRAMs, some errors favor one direction or the other. If this bias is large enough, the memory controller can use it to help error correction. MUSE-ECC’s residue codes [28] implement error correction based on integer arithmetic instead of $\mathbb{F}_{2^{n}}$ arithmetic, enabling asymmetric codes. This includes an $(80,67)$ code (measured in bits) that can correct any $1\!\!\!\to\!\!\!0$ 8-bit symbol error. This exceeds the Komamiya-Singleton bound, which applies only to symmtric channels.
But asymmetric codes are not as applicable in modern DRAMs. Energy-saving changes reduce the directional bias of errors, and DDR5’s on-die ECC can miscorrect, adding an extra bit-flip in either direction. MUSE-ECC might be adaptable to this model, but this remains to be seen. We would also like further reliability features such as erasure decoding.
MUSE-ECC also allows symmetric codes with shorter symbols than Reed-Solomon would permit, but since our work focuses on reliability we prefer longer symbols .
V-B Comparison to other large-codeword techniques
@@
DUO ECC goes further, by repurposing the on-die ECC data to provide better controller-level error correction with a long codeword, at the cost of slightly longer transfers, possibly enabling probabilistic CC or even DCC for DIMM configurations that would otherwise not support them. For example, DUO suggests DDR4 ×$\times 4$ and DDR4 ×$\times 8$ modes with 18 and 9 devices, 34 and 68 bits per device, respectively, both with a total of $34\cdot 18=612$-bit codewords. These use a $\mathrm{RS}(2^{8};76,64)$ code which can correct up to 6 byte errors. Or they can probabilistically correct a single-device error for DDR5, or even a double-device error for DDR4, by attempting decoding with each device (or pair of devices) erased.
As with Bamboo codes, URS codes can complement DUO, but not as effectively because there are not a whole number of symbols per device, which necessitates shortening the codes [^8] Unless 17-bit symbols are used, but this is worse unless significant metadata is also desired, because it would have a $31\cdot 17=527$-bit payload.. For example, we could use $\mathbb{F}_{2^{9}}$ and divide the 100 extra bits into 1 bit of metadata and 11 nine-bit parity symbols. This would give an $\mathrm{RS}(2^{9};72,61)$ code with 4 or 8 symbols per device for DDR4 or DDR5, shortened from 72 or 36 bits to 68 or 34 bits per device, respectively. In DDR4 this corrects a single-device plus one error efficiently and reliably; a double-device error with failure probability about $2^{-18}\approx 3.81\mathrm{e}{-6}$; or a single-device plus two-symbol error with failure probability about $2^{-9}\approx 0.2\%$. With DDR5, it can correct a single-device failure except with probability about $2^{-27}\approx 7.45\mathrm{e}{-9}$. These calculations are all based on analytically solving the key equation rather than iterating over each device. We could fall back to the iterative solution if the analytic solution fails, which would be a nearly identical solution to DUO but with slightly higher efficiency due to the sparse erasure locator polynomial.
@@ VI Conclusion and future work
VI-A Acknowledgments
Many thanks to Tanj Bennett for helpful discussion, and to anonymous reviewers for their feedback.
VI-B Intellectual property disclosure